[Brahms-l] Shift report 20050529 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 29 2005 - 15:56:36 EDT
Shifter leader was HHD
Shifters were 


1. One stable store 0800-1100 and a rather unstable one 1300-1500. 

2. Data runs taken:
   FS@4A3450,MRS@90A1050: 15053,15054,15055,15056,15057
   FS@4A3450,MRS@90A700: 15060,15061,15062
   Non-TPC pedestal run: 15058
   TPC pedestal run: 15059
3. BB voltages are acting weird. Some of them drops suddenly which can         be fixed temporarily by turning on the voltages (in BRAHVO) several times ignoring error messages. 

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 29 15:56:48 2005

This archive was generated by hypermail 2.1.8 : Sun May 29 2005 - 15:56:56 EDT